Circular No. 7546

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2000 WM_{1} (LINEAR) An apparently asteroidal object with unusual motion reported by the

LINEAR team on Dec. 16 was posted on the NEO Confirmation Page. Subsequent astrometry permitted a linkage to another set of observations by LINEAR on Nov. 16 and 18, designated 2000 WM_{1} on MPS 22800. The first LINEAR observations on Nov. 16 and Dec. 16 are given below. An observation of 2000 WM_{1} by T. B. Spahr (Smithsonian Astrophysical Observatory 1.2-m reflector at Mt. Hopkins) on Dec. 20.148 UT shows the object to have a 10′′ coma and a broad, faint tail some 10′′–20′′ long in p.a. 45^{o}. The full set of observations, together with the parabolic orbital elements given below and an ephemeris that indicates the comet may reach naked-eye brightness late next year, appears on MPEC 2000-Y20.

2000 UT  α_{2000}  δ_{2000}  m_{2}
Nov. 16.14461  0^{h}08^{m}34.^{s}62  +59^{o}47′36′′.0  18.0
Dec. 16.06960  23 58 17.08  +54 29 11.3  17.8
T = 2002 Jan. 22.7781 TT  ω = 276.^{o}8165
Ω = 237.8924 }2000.0
q = 0.554808 AU  i = 72.5589

SUPERNOVA 2000fn IN NGC 2526 T. Puckett, Mountain Town, GA, provides his astrometry and pho-

tometry of an apparent supernova (mag ∼ 18.8) discovered by B. Holmes, Charleston, IL, on an unfiltered CCD frame taken with a 0.40-m reflector on Dec. 17.20 UT: α = 8^{h}06^{m}57^{s}.79, δ = +8^{o}00′19′′.3 (equinox 2000.0), which is 12′′.7 west and 4′′.9 north of the center of NGC 2526. The new object was also present (mag 18.0) on an unfiltered CCD frame (limiting mag 19.0) taken by Holmes on Nov. 27.20, and (via W. Li, University of California) on images taken with the Katzman Automatic Imaging Telescope on Nov. 19 and Dec. 3. SN 2000fn does not appear on Palomar Sky Survey images (limiting mag ∼ 21.0) taken on 1997 Nov. 7 and 1988 Dec. 11.

COMET C/2000 W1 (UTSUNOMIYA-JONES) Visual m_{1} estimates (B = binoculars): Dec. 5.93 UT, 7.0 (J. G. S.

Aguiar, Campinas, Brazil, 11×80 B); 6.97, 6.5 (A. Amorim, Florianopolis, Brazil, 7×50 B; moonlight); 13.52, 6.7 (A. Pearce, Nedlands, W. Australia, 20×80 B); 16.37, 5.5 (S. Yoshida, Ibaraki, Japan, 0.25-m reflector).

2000 December 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green